please help!!! Angle B (sry don't know what symbol to use for angle) is supplementary to Angle A determine m(Angle)B Angle A = 42 degrees Thanks for helping.
Answer by stanbon(48529)

Angle B (sry don't know what symbol to use for angle) is supplementary to Angle A determine m(Angle)B Angle A = 42 degrees
---------
A + B = 180 degrees
42 + B = 180
B = 180-42 = 138 degrees
==========================
Cheers,
Stan H.